OpenStudy (callisto):

C and D are two points on AB and OB respectively such that AC:AB=1:3 and OD:OB=2:5. Let OA =a and OB =b. (a) express OC and AD in terms of a and b (b)AD intersect CO at E such that AE:AD = 1:m and OE:OC = 1:n. Find the values of m and n My problem is that how to set the 2 equations for (b) so that i can solve for m and n?

OpenStudy (callisto):

Here's the diagram. It is a vector question..

OpenStudy (mani_jha):

Is AD a median and OC an altitude?

OpenStudy (callisto):

OD:OB = 2:5, so i don't think AD is the median

OpenStudy (callisto):

actually, both are not mentioned in the question

OpenStudy (mani_jha):

I can't write the vector notation(arrow on top), so I writing it the regular way. OA=a OD=2b/5 AD=OA+OD(Vectorically)

OpenStudy (callisto):

my mistake sorry

OpenStudy (mani_jha):

Oh yes that it should be. It was my mistake.

OpenStudy (callisto):

but what's next?

OpenStudy (mani_jha):

Do you need AD in vector form? Then just substitute the values of AO and OD...

OpenStudy (callisto):

i've solved part a already. My problem is that how to set the 2 equations for (b) so that i can solve for m and n?

OpenStudy (mani_jha):

Well, I am working on that part too. Have you tried: AD=AE+ED OC=OE+EC We know the values of AD and OC. Now find AE and ED in terms of AD, I found it: AE=AD/m ED=(m-1)AD/m Similarly, do so for OE and EC in terms of OC. Then add the above two equations and see what you get. I will post as soon as I get an update on this.

OpenStudy (callisto):

i'm working on OE +ED = OD and AC+CE=AE but seems to be very complicated

OpenStudy (callisto):

if you add AE=AD/m and ED=(m-1)AD/m together, probably, you'll get AD=AD

OpenStudy (mani_jha):

try to add AE and AC instead. and also OE and OD

OpenStudy (callisto):

then, that's quite similar to my method

OpenStudy (callisto):

My workings for this question is on the right hand side.. Can you check it for me?

OpenStudy (mani_jha):

It's correct. Good work.

OpenStudy (callisto):

Thank you